OpenStudy (harleygirl18):

What is the difference -1/10 -8/10? I think the answer is a but I wanna be sure A.-9/10 B.-7/10 C.7/10 D.9/10

OpenStudy (calculusxy):

Since the denominator is the same we can just subtract the numerators: \[\large \frac{ -1 - 8 }{ 10 } = \frac{-9}{10}\]

OpenStudy (harleygirl18):

Thank you sooo much calculusxy